OpenStudy (anonymous):

Boat crew rowed 10.5 mi downstream, with the current in 1.5 mi. hr. the return trip upstream against the current covered the same distance but took 3.5 hrs. find the crew rowing rate in still water and the rate of the current.

OpenStudy (turingtest):

is the distance 10.5mi and the first time 1.5 hr (you wrote mi. hr)?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

v=d/t let x=speed of the boat in still water y=speed of the current vt=(x+y)(1.5)=10.5--->x+y=7 (1) on the return trip: vt=(x-y)(3.5)=10.5---> x-y=3 (2) add (1) to (2): 2x=10 x=5 use this to solve for y: y=7-x=2 so x=5 y=2

OpenStudy (anonymous):

thank you!